OpenStudy (anonymous):

write an equation in point slope form of the line through point J (4,1) with slope -4

OpenStudy (anonymous):

HELP SOMEONE PLEASE

undeadknight26 (undeadknight26):

y = mx + b?

OpenStudy (anonymous):

Which one is right? @happytales @undeadknight26

undeadknight26 (undeadknight26):

i gave u the base formula....happy gave you it with everything in it...so technically they are both right XD

OpenStudy (anonymous):

So happy y=-4x+1 is right, correct?

OpenStudy (anonymous):

uhm no?

OpenStudy (e.mccormick):

Point slope is: \(y-y_1=m(x-x_1)\) Slope Intercept is \(y=mx+b\)

OpenStudy (e.mccormick):

The point in point slope form is given to you as something like: \((x_1, y_1) = (5, 3)\). That is just a random example. Your problem has (4, 1) so put that into the first equation above.